OpenStudy (anonymous):

Marla bought one necklace and two bracelets for $67.50. Use the following information to determine the amount she paid for each bracelet: •The price of the necklace was $27. •The bracelets were both the same price. Enter your numerical answer only in the box; do not include the dollar sign.

OpenStudy (anonymous):

67.50-27?

OpenStudy (texaschic101):

1n + 2b = 67.50 n = 27 now we sub 27 in for n 27 + 2b = 67.50 2b = 67.50 - 27 2b = 40.50 b = 40.50/2 b = 20.25 check.. 1n + 2b = 67.50 1(27) + 2(20.25) = 67.50 27 + 40.50 = 67.50 67.50 = 67.50 (correct) so the bracelets (b) cost 20.25 each

OpenStudy (texaschic101):

do you have any questions at all ?

OpenStudy (anonymous):

am taking note of your solution..its clear. thanks

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

Which of the following expressions are equivalent to −3.5(2 − 3n) − 2.5n?

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

@texaschic101 can you help me with -3.5(2-3n)-2.5n

OpenStudy (anonymous):

-7+8

OpenStudy (texaschic101):

-3.5(2 - 3n) - 2.5n --- distribute through the parenthesis -7 + 10.5n - 2.5n --- combine like terms -7 + 8n

OpenStudy (anonymous):

i did that they say its wrong

OpenStudy (anonymous):

A. −7 − 8n B. −7 + 8n C. −7 − 13n D. −7 − n(10.5 − 2.5) E. −7 + n(10.5 − 2.5)

OpenStudy (texaschic101):

answer is B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

can you assit last one

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

Solve each equation. Then, enter a number in each box to make this statement true

OpenStudy (anonymous):

3/4(x-12)=12

OpenStudy (anonymous):

3/4y-12=12

OpenStudy (texaschic101):

is there parenthesis in the second equation ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

3/4(x - 12) = 12 3/4x - 9 = 12 3/4x = 12 + 9 3/4x = 21 x = 21 / (3/4) x = 21 * 4/3 x = 84/3 = 28 3/4y - 12 = 12 3/4y = 12 + 12 3/4y = 24 y = 24/(3/4) y = 24 * 4/3 y = 96/3 = 32 Now what statement do I have to make true ?

OpenStudy (anonymous):

x and y

OpenStudy (anonymous):

all i have to do is substitute x and y with answer ?

OpenStudy (texaschic101):

I solved both equations...the 1st one is : x = 28 and the 2nd one is y = 32

OpenStudy (texaschic101):

those would make your equations true

OpenStudy (anonymous):

thank you very much

OpenStudy (texaschic101):

glad to help :)